OpenStudy (anonymous):

Which of the following is the general form of the equation 5x2 - 20x = 15? What is the vertex of the parabola whose equation is y = (x + 1)2 + 3? Describe the translation of the graph of y = x2 that results in the graph of y = (x - 3)2.

OpenStudy (anonymous):

Which of the following is the general form of the equation 5x2 - 20x = 15? the one that looks like \(5x^2-20x-15=0\)

OpenStudy (anonymous):

general form always = 0 right

OpenStudy (anonymous):

right

OpenStudy (anonymous):

What is the vertex of the parabola whose equation is \(y = (x + 1)^2 + 3\) it is in vertex form \(y=(x-h)^2+k\) with \(h=-1,k=3\) making the vertex \((-1,3)\)

OpenStudy (anonymous):

isnt the last one right 3 units?

OpenStudy (anonymous):

Describe the translation of the graph of \(y = x^2\) that results in the graph of \(y = (x - 3)^2\)

OpenStudy (anonymous):

yes it is

OpenStudy (anonymous):

thanks dad!

OpenStudy (anonymous):

https://www.youtube.com/watch?v=wT-ARHa3Yr8

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

lol! im gonna go play league of legends, i hope youll be here to help me when i get back

OpenStudy (anonymous):

i gotta go take my geritol and put on a van morrison lp and go to bed

OpenStudy (anonymous):

oh okay good night dad, im gonna go rave and be a groupie

OpenStudy (anonymous):

k see you in the morning

OpenStudy (anonymous):

leave the back door unlocked i forget how keys work when its late

OpenStudy (anonymous):

i'll leave the door open if you can find it